# The moduli spaces of left-invariant statistical structures on Lie groups

### v1

@@ 6.3. Amari–Chentsov $\alpha$-connection on the Takano Gaussian space
One sees that the Lie group $G_{\mathbb{R}\mathrm{H}^{m+1}}$ is isomorphic to $\mathrm{Aff}^{d+}(m,\mathbb{R})$. Moreover, $\mathrm{Aff}^{d+}(m,\mathbb{R})$ has a simply-transitive action on $\mathcal{N}_{T}^{m}$ under which the Fisher metric $g^{F}$ and the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ on $\mathcal{N}_{T}^{m}$ are invariant, see Example 3.1.2. Thus, by Theorem 6.1.1 and Proposition 2.4.5, we obtain the claim. ∎
@@
Theorem 6.3.1 provides a characterization similar to the above Proposition 6.3.2 for the Amari–Chentsov $\alpha$-connection on $(\mathcal{N}_{T}^{m},\mathbb{R}^{m},\mathbf{p}_{T})$.
Let us consider the case where $m=2$. In [13], the Lie group $\mathrm{Aff}^{d+}(1,\mathbb{R})=\mathbb{R}_{>0}\ltimes\mathbb{R}$ is called the Lie group of (univariate) normal distributions. Note that $\mathrm{Aff}^{d+}(1,\mathbb{R})\cong G_{\mathbb{R}\mathrm{H}^{2}}$. By Theorem 6.1.1, for the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(\mathrm{Aff}^{d+}(1,\mathbb{R}))$, the following theorem holds.
Theorem 6.3.3.
@@
$S:=\left\{(a,b,c,d)\mid b>0\right\}\ \cup\ \left\{(a,0,c,d)\mid d\geq 0\right\}\subset S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{2}}^{\ast}).$
@@ 7. The case of $H^{3}\times\mathbb{R}^{n-3}$
In this section, we consider the product Lie group of the $3$-dimensional Heisenberg group $H^{3}$ and the commutative Lie group $\mathbb{R}^{n-3}$, where $n\geq 3$. Then $\mathfrak{PM}(H^{3}\times\mathbb{R}^{n-3})\cong\{\ast\}$ (see Proposition 4.1.2).
@@ 7.1. Main theorem for $H^{3}\times\mathbb{R}^{n}$
As supplementary remarks to Theorem 7.1.1, we will also discuss the following points:
Proposition 7.1.4.
@@
In this subsection, we give proofs of Theorem 7.1.1 and Proposition 7.1.4.
@@
Since $\mathfrak{PM}(H^{3}\times\mathbb{R}^{n-3})\cong\{\ast\}$, by applying Proposition 4.3.1 and 4.3.2 (2), we have
@@
Therefore Theorem 7.1.1 and Proposition 7.1.4 follows from the three propositions below.
Proposition 7.2.1.
@@
First, we shall give a proof of Proposition 7.2.1:
Proof of Proposition 7.2.1.
@@
In order to prove Proposition 7.2.2, we prepare the following lemma:
Lemma 7.2.4.
@@
Proof of Lemma 7.2.4.
By definition, the structure constants of $\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$ with respect to the canonical basis are given by $a_{12}^{3}=1$, $a_{21}^{3}=-1$, with all other $a_{ij}^{k}$ equal to zero. Then, by applying Proposition 3.3.1, the generalized Christoffel symbols can be computed as
@@
with all other $\Gamma_{ij}^{k}$ equal to zero. The claims of Lemma 7.2.4 can be obtained by applying Proposition 3.3.2. ∎
Let us give a proof of Proposition 7.2.2.
Proof of Proposition 7.2.2.
@@
and $C=\sum_{u,v,s}C_{uvs}(x_{u}\otimes x_{v}\otimes x_{s})$. Then by Lemma 7.2.4, the following claims hold for each $i,j\geq 4$:
@@
Then, it immediately follows from Lemma 7.2.4 that $\nabla^{g}C=0$. ∎
Finally, we shall prove Proposition 7.2.3:
Proof of Proposition 7.2.3.
@@
Recall that $R^{0}_{g}=\frac{3}{4}\omega_{12}\odot\omega_{12}-\frac{1}{4}\omega_{13}\odot\omega_{13}-\frac{1}{4}\omega_{23}\odot\omega_{23}$ (see [27]). Thus by applying Proposition 2.3.2, our goal is to show that
@@ Appendix A Remarks on topologies of equivariant fiber bundles
In this section, we recall a theorem (Theorem A.0.1) concerning topologies of equivariant fiber bundles, which is applied in Section 4.3.
@@
Note that for each $v\in F$ and each $g\in O$, the equalities $\Phi(e_{G},v)=v$ and $(g.x,v)=\eta(g.\alpha)$ holds where we define $\alpha$ the unique element in $E_{x}$ with $\eta_{F}(\alpha)=\Phi(g,v)\in F$. By Lemma A.0.2 below, one can find an open neighborhood $O^{\prime}$ of $e_{G}$ in $O$ such that $\Phi(O^{\prime}\times K)\subset V_{x}^{\prime}$. For such $O^{\prime}$, one can see that
@@ Appendix B Remarks on moduli of geometric structures on homogeneous spaces
Proof of Theorem B.0.3 can be given similarly to the proof of Theorem B.0.2.
Note that Proposition 4.2.2 is a special case of Theorem B.0.3 when $H$ is trivial and $M=G$.
@@
Remark C.0.2.
We believe that $\Xi$ is a smooth submanifold of the vector space $S^{3}((\mathbb{R}^{n})^{\ast})$ and that the natural $O(n)$-action on $\Xi$ is polar with a section $V$. If this is indeed the case, then Theorem C.0.1 would follow directly from the general theory of polar actions (cf. [11]). Although we have attempted to verify this claim, a proof has not yet been obtained.
@@
A key part of Theorem C.0.1 is the following:
Proposition C.0.3.
@@
To prove Proposition C.0.3, we give two lemmas below:
Lemma C.0.4.
@@
$\mathrm{supp}(\lambda):=\{i\in\{1,\dots,n\}\mid\lambda_{i}\neq 0\}$
@@
Proof of Proposition C.0.3.
For each $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$, we write $C_{\lambda}=\sum_{i}\lambda_{i}x_{i}^{3}\in V$. Fix $\lambda,\mu\in\mathbb{R}^{n}$ and $k\in O(n)$ with $k.C_{\lambda}=C_{\mu}$. Then our goal is to find $k^{\prime}\in W$ such that $k^{\prime}.C_{\lambda}=C_{\mu}$. By Lemma C.0.4, $\Omega_{\lambda}\cap k^{-1}\Omega_{\mu}$ is open dense in $\mathbb{R}^{n}$, and in particular non-empty. Fix $u=\sum_{i}a_{i}e_{i}\in\Omega_{\lambda}\cap k^{-1}\Omega_{\mu}$. We write $ku=\sum_{i}b_{i}e_{i}\in\Omega_{\mu}$. Let us define the linear endomorphisms $\widetilde{K}^{C_{\lambda}}_{u}$ and $\widetilde{K}^{C_{\mu}}_{ku}$ on $\mathbb{R}^{n}$ as in Lemma C.0.5. Then by the definition, we have $\widetilde{K}^{C_{\lambda}}_{u}(e_{i})=a_{i}\lambda_{i}e_{i}$ and $\widetilde{K}^{C_{\mu}}_{ku}(e_{i})=b_{i}\mu_{i}e_{i}$ for each $i$. Since $u\in\Omega_{\lambda}$, we see that the family
@@
gives the eigenspace decomposition of $\mathbb{R}^{n}$ by the operator $\widetilde{K}^{C_{\mu}}_{ku}$ such that the eigenvalue of $\widetilde{K}^{C_{\mu}}_{ku}$ on $\mathrm{Span}\{e_{i}\mid\mu_{i}=0\}$ is zero. By Lemma C.0.5,
@@
Proof of Theorem C.0.1.
The continuous map $\psi:V_{+}\rightarrow O(n)\backslash\Xi$ is surjective because any $W$-orbit in $V$ intersects $V_{+}$ and $W\subset O(n)$. The injectivity of $\psi$ follows from Proposition C.0.3 as below: Let us take any two elements $P,Q\in V_{+}$ such that $O(n).P=\psi(P)=\psi(Q)=O(n).Q$. Then by Proposition C.0.3, $W.P=W.Q$. Since $V_{+}$ gives a complete representative of $W$-action on $V$, we have $P=Q$. This proves that the map $\psi$ is injective.
@@
  • [7] Y. Fujitani. Information geometry of warped product spaces. Inf. Geom., 6:127–155, 2023. https://doi.org/10.1007/s41884-022-00091-9
@@
  • [11] C. Gorodski. Topics in polar actions. arXiv, arXiv:2208.03577, 2025. https://doi.org/10.48550/arXiv.2208.03577
@@
  • [14] J. Inoguchi and Y. Ohno. Homogeneous statistical manifolds. Inf. Geom., 2025. https://doi.org/10.1007/s41884-025-00172-5
@@
  • [16] S.-P. Kobayashi and Y. Ohno. On a constant curvature statistical manifold. Inf. Geom., 5:31–46, 2022. https://doi.org/10.1007/s41884-022-00065-x
  • [17] S.-P. Kobayashi and Y. Ohno. A characterization of the alpha-connections on the statistical manifold of multivariate normal distributions. Osaka J. Math., 62(2):22, 2025. https://doi.org/10.18910/101132
  • [18] H. Kodama, A. Takahara, and H. Tamaru. The space of left-invariant metrics on a Lie group up to isometry and scaling. Manuscripta Math., 135:229–243, 2011. https://doi.org/10.1007/s00229-010-0419-4
@@
  • [22] T. Kurose. Dual connections and affine geometry. Math. Z., 203:115–121, 1990. https://doi.org/10.1007/BF02570725
@@
  • [28] K. Nomizu and T. Sasaki. Affine differential geometry: geometry of affine immersions. Cambridge Univ. Press, 1994.
@@
  • [32] S. M. Salamon. Complex structures on nilpotent Lie algebras. J. Pure Appl. Algebra, 157(2–3):311–333, 2001. https://doi.org/10.1016/S0022-4049(00)00033-5
  • [33] H. Shima. Homogeneous hessian manifolds. Ann. Inst. Fourier (Grenoble), 30:91–128, 1980. https://doi.org/10.5802/aif.794
@@
  • [35] H. Shima. The Geometry of Hessian Structures. World Sci., 2007. https://doi.org/10.1142/6241
@@
  • [37] Y. Taketomi. A maximal element of a moduli space of Riemannian metrics. arXiv, arXiv:2210.01483, 2022. https://doi.org/10.48550/arXiv.2210.01483

### v3

@@ 6.3. Amari–Chentsov $\alpha$-connection on the Takano Gaussian space
One sees that the Lie group $G_{\mathbb{R}\mathrm{H}^{m+1}}$ is isomorphic to $\mathrm{Aff}^{d+}(m,\mathbb{R})$. Moreover, $\mathrm{Aff}^{d+}(m,\mathbb{R})$ has a simply-transitive action on $\mathcal{N}_{T}^{m}$ under which the Fisher metric $g^{F}$ and the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ on $\mathcal{N}_{T}^{m}$ are invariant, see Example 3.2. Thus, by Theorem 6.1 and Proposition 2.14, we obtain the claim. ∎
@@
Theorem 6.8 provides a characterization similar to the above Proposition 6.9 for the Amari–Chentsov $\alpha$-connection on $(\mathcal{N}_{T}^{m},\mathbb{R}^{m},\mathbf{p}_{T})$.
Let us consider the case where $m=2$. In [12], the Lie group $\mathrm{Aff}^{d+}(1,\mathbb{R})=\mathbb{R}_{>0}\ltimes\mathbb{R}$ is called the Lie group of (univariate) normal distributions. Note that $\mathrm{Aff}^{d+}(1,\mathbb{R})\cong G_{\mathbb{R}\mathrm{H}^{2}}$. By Theorem 6.1, for the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(\mathrm{Aff}^{d+}(1,\mathbb{R}))$, the following theorem holds.
Theorem 6.10.
@@
$S:=\left\{(a,b,c,d)~\middle|~b>0\right\}\ \cup\ \left\{(a,0,c,d)~\middle|~d\geq 0\right\}\subset S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{2}}^{\ast}).$
@@ 7. The case of $H^{3}\times\mathbb{R}^{n-3}$
In this section, we consider the product Lie group of the $3$-dimensional Heisenberg group $H^{3}$ and the commutative Lie group $\mathbb{R}^{n-3}$, where $n\geq 3$. Then $\mathfrak{PM}(H^{3}\times\mathbb{R}^{n-3})\cong\{\ast\}$ (see Proposition 4.2).
@@ 7.1. Main theorem for $H^{3}\times\mathbb{R}^{n}$
As supplementary remarks to Theorem 7.1, we will also discuss the following points:
Proposition 7.4.
@@
In this subsection, we give proofs of Theorem 7.1 and Proposition 7.4.
@@
Since $\mathfrak{PM}(H^{3}\times\mathbb{R}^{n-3})\cong\{\ast\}$, by applying Propositions 4.6 and 4.7 (2), we have
@@
Therefore Theorem 7.1 and Proposition 7.4 follows from the three propositions below.
Proposition 7.5.
@@
First, we shall give a proof of Proposition 7.5:
Proof of Proposition 7.5.
@@
In order to prove Proposition 7.6, we prepare the following lemma:
Lemma 7.8.
@@
Proof of Lemma 7.8.
By definition, the structure constants of $\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$ with respect to the canonical basis are given by $a_{12}^{3}=1$, $a_{21}^{3}=-1$, with all other $a_{ij}^{k}$ equal to zero. Then, by applying Proposition 3.4, the generalized Christoffel symbols can be computed as
@@
with all other $\Gamma_{ij}^{k}$ equal to zero. The claims of Lemma 7.8 can be obtained by applying Proposition 3.5. ∎
Let us give a proof of Proposition 7.6.
Proof of Proposition 7.6.
@@
and $C=\sum_{u,v,s}C_{uvs}(x_{u}\otimes x_{v}\otimes x_{s})$. Then by Lemma 7.8, the following claims hold for each $i,j\geq 4$:
@@
Then, it immediately follows from Lemma 7.8 that $\nabla^{g}C=0$. ∎
Finally, we shall prove Proposition 7.7:
Proof of Proposition 7.7.
@@
Recall that $R^{0}_{g}=\frac{3}{4}\omega_{12}\odot\omega_{12}-\frac{1}{4}\omega_{13}\odot\omega_{13}-\frac{1}{4}\omega_{23}\odot\omega_{23}$ (see [26]). Thus by applying Proposition 2.7, our goal is to show that
@@ Appendix A Remarks on topologies of equivariant fiber bundles
In this section, we recall a theorem (Theorem A.1) concerning topologies of equivariant fiber bundles, which is applied in Section 4.3.
@@
Note that for each $v\in F$ and each $g\in O$, the equalities $\Phi(e_{G},v)=v$ and $(g.x,v)=\eta(g.\alpha)$ holds where we define $\alpha$ the unique element in $E_{x}$ with $\eta_{F}(\alpha)=\Phi(g,v)\in F$. By Lemma A.2 below, one can find an open neighborhood $O^{\prime}$ of $e_{G}$ in $O$ such that $\Phi(O^{\prime}\times K)\subset V_{x}^{\prime}$. For such $O^{\prime}$, one can see that
@@ Appendix B Remarks on moduli of geometric structures on homogeneous spaces
Proof of Theorem B.3 can be given similarly to the proof of Theorem B.2.
Note that Proposition 4.4 is a special case of Theorem B.3 when $H$ is trivial and $M=G$.
@@
Remark C.2.
We believe that $\Xi$ is a smooth submanifold of the vector space $S^{3}((\mathbb{R}^{n})^{\ast})$ and that the natural $O(n)$-action on $\Xi$ is polar with a section $V$. If this is indeed the case, then Theorem C.1 would follow directly from the general theory of polar actions (cf. Gorodski (2022), arXiv:2208.03577). Although we have attempted to verify this claim, a proof has not yet been obtained.
@@
A key part of Theorem C.1 is the following:
Proposition C.3.
@@
To prove Proposition C.3, we give two lemmas below:
Lemma C.4.
@@
$\mathrm{supp}(\lambda):=\left\{i\in\{1,\dots,n\}~\middle|~\lambda_{i}\neq 0\right\}$
@@
Proof of Proposition C.3.
For each $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$, we write $C_{\lambda}=\sum_{i}\lambda_{i}x_{i}^{3}\in V$. Fix $\lambda,\mu\in\mathbb{R}^{n}$ and $k\in O(n)$ with $k.C_{\lambda}=C_{\mu}$. Then our goal is to find $k^{\prime}\in W$ such that $k^{\prime}.C_{\lambda}=C_{\mu}$. By Lemma C.4, $\Omega_{\lambda}\cap k^{-1}\Omega_{\mu}$ is open dense in $\mathbb{R}^{n}$, and in particular non-empty. Fix $u=\sum_{i}a_{i}e_{i}\in\Omega_{\lambda}\cap k^{-1}\Omega_{\mu}$. We write $ku=\sum_{i}b_{i}e_{i}\in\Omega_{\mu}$. Let us define the linear endomorphisms $\widetilde{K}^{C_{\lambda}}_{u}$ and $\widetilde{K}^{C_{\mu}}_{ku}$ on $\mathbb{R}^{n}$ as in Lemma C.5. Then by the definition, we have $\widetilde{K}^{C_{\lambda}}_{u}(e_{i})=a_{i}\lambda_{i}e_{i}$ and $\widetilde{K}^{C_{\mu}}_{ku}(e_{i})=b_{i}\mu_{i}e_{i}$ for each $i$. Since $u\in\Omega_{\lambda}$, we see that the family
@@
gives the eigenspace decomposition of $\mathbb{R}^{n}$ by the operator $\widetilde{K}^{C_{\mu}}_{ku}$ such that the eigenvalue of $\widetilde{K}^{C_{\mu}}_{ku}$ on $\mathrm{Span}\{e_{i}\mid\mu_{i}=0\}$ is zero. By Lemma C.5,
@@
Proof of Theorem C.1.
The continuous map $\psi:V_{+}\rightarrow O(n)\backslash\Xi$ is surjective because any $W$-orbit in $V$ intersects $V_{+}$ and $W\subset O(n)$. The injectivity of $\psi$ follows from Proposition C.3 as below: Let us take any two elements $P,Q\in V_{+}$ such that $O(n).P=\psi(P)=\psi(Q)=O(n).Q$. Then by Proposition C.3, $W.P=W.Q$. Since $V_{+}$ gives a complete representative of $W$-action on $V$, we have $P=Q$. This proves that the map $\psi$ is injective.
@@
  • [7] Y. Fujitani. Information geometry of warped product spaces. Inf. Geom., 6(1):127–155, 2023. https://doi.org/10.1007/s41884-022-00091-9
@@
  • [13] J. Inoguchi and Y. Ohno. Homogeneous statistical manifolds. Inf. Geom., 8(2):285–341, 2025. https://doi.org/10.1007/s41884-025-00172-5
@@
  • [15] S.-P. Kobayashi and Y. Ohno. On a constant curvature statistical manifold. Inf. Geom., 5(1):31–46, 2022. https://doi.org/10.1007/s41884-022-00065-x
  • [16] S.-P. Kobayashi and Y. Ohno. A characterization of the alpha-connections on the statistical manifold of multivariate normal distributions. Osaka J. Math., 62(2):329–349, 2025. https://doi.org/10.18910/101132
  • [17] H. Kodama, A. Takahara, and H. Tamaru. The space of left-invariant metrics on a Lie group up to isometry and scaling. Manuscripta Math., 135(1-2):229–243, 2011. https://doi.org/10.1007/s00229-010-0419-4
@@
  • [21] T. Kurose. Dual connections and affine geometry. Math. Z., 203(1):115–121, 1990. https://doi.org/10.1007/BF02570725
@@
  • [27] K. Nomizu and T. Sasaki. Affine Differential Geometry, vol. 111 of Cambridge Tracts in Mathematics. Cambridge Univ. Press, Cambridge, 1994.
@@
  • [31] S. M. Salamon. Complex structures on nilpotent Lie algebras. J. Pure Appl. Algebra, 157(2-3):311–333, 2001. https://doi.org/10.1016/S0022-4049(00)00033-5
  • [32] H. Shima. Homogeneous Hessian manifolds. Ann. Inst. Fourier (Grenoble), 30(3):91–128, 1980. https://doi.org/10.5802/aif.794
@@
  • [34] H. Shima. The Geometry of Hessian Structures. World Sci., Singapore, 2007. https://doi.org/10.1142/6241